# Raindrop GS: A Benchmark for 3D Gaussian Splatting under Raindrop Conditions

### v1

@@ 3.3 Raindrop-aware 3DGS Evaluation
With the estimated camera poses and initialized point cloud, we proceed to evaluate four representative 3DGS methods: 3DGS Kerbl et al. (2023), WeatherGS Qian et al. (2024), GS-W Zhang et al. (2024), and 3DGS-MCMC Kheradmand et al. (2024).
Among these, 3DGS Kerbl et al. (2023) serves as a standard baseline for 3D Gaussian splatting. WeatherGS Qian et al. (2024) incorporates single-image raindrop removal, so we omit the explicit raindrop removal step in its data processing pipeline. GS-W Zhang et al. (2024) is specifically designed for challenging conditions and unconstrained image collections, making it more robust to inconsistent multi-view inputs. 3DGS-MCMC Kheradmand et al. (2024), on the other hand, does not rely on accurate point cloud initialization. Each of the aforementioned methods has its own advantages, making their evaluation in our benchmark both meaningful and insightful.
Figure 4: Qualitative comparison across different data preprocessing strategies: Uformer and Restormer exhibit comparable raindrop removal capability, while COLMAP, when successful in estimating camera poses, reconstructs finer details than VGGT.
@@ 4.1 Implementation Details
We standardize all scene images to a resolution of 1024 * 576 for uniform comparison. During camera-pose estimation, point-cloud initialization, and raindrop removal, we follow the settings and recommendations provided by the respective authors to ensure each method operates at its optimal performance. For all 3DGS methods, we adhere to their official optimization settings, we run 30,000 iterations for 3DGS, WeatherGS, and 3DGS-MCMC, and 70,000 iterations for GS-W. For 3DGS-MCMC, we set the initial point cloud size to 100,000, based on the number of points that VGGT can initialize in our dataset.
The evaluation employs five metrics: PSNR, SSIM Wang et al. (2004), LPIPS Zhang et al. (2018), AUC@30 Wang et al. (2023), and 3D Points. Among them, PSNR, SSIM, and LPIPS assess image deraining and 3DGS reconstruction performance; AUC@30 evaluates camera pose estimation accuracy; and 3D Points measures the number of point cloud initialization. All models are implemented in PyTorch and trained concurrently on 8 NVIDIA RTX 3090 GPUs.
@@ 4.2 Quantitative Comparison
Table 4 compares the impact of background-focused (BG-focused) and raindrop-focused (RD-focused) captures on 3DGS performance using VGGT Wang et al. (2025). For the original 3DGS method, raindrop-focused images exhibit a 4 dB drop compared to background-focused images, due to background blur and light refraction. VGGT processes all scenes but generates a point cloud with 0 points for raindrop-focused images, for which we use a randomly initialized point cloud.
Table 2 and Table 5 compares the performance of VGGT Wang et al. (2025) and COLMAP Schonberger & Frahm (2016b). For camera pose estimation, we use VGGT and COLMAP to estimate poses on the ground-truth images and compare these estimates with the poses obtained from the corresponding images. VGGT yields more accurate camera pose estimates; both methods accurately recover poses for background-focused images but exhibit substantial performance degradation on raindrop-focused images. In terms of point cloud initialization, VGGT outperforms COLMAP for background point clouds, yet it fails to initialize the raindrop-focused dataset. However, when COLMAP successfully produces camera poses and an initialized point cloud, the original 3DGS method achieves the best performance with Uformer Wang et al. (2022).
Table 3 reports the performance of different deraining/restoration methods on raindrop-affected scenes. Uformer Wang et al. (2022) and Restormer Zamir et al. (2022) show comparable results, with Uformer Wang et al. (2022) slightly outperforming Restormer Zamir et al. (2022) in reconstruction metrics, while IDT Xiao et al. (2022) lags substantially behind the other two methods.
Table 2: Comparison of camera pose estimation (AUC@30 Wang et al. (2023)) and the number of points in the point cloud between VGGT and COLMAP on background- and raindrop-focused datasets.
@@
3D Points 69401.11 0 5476.89 302.50
For 3DGS methods, GS-W Zhang et al. (2024) with VGGT Wang et al. (2025) and Uformer Wang et al. (2022) preprocessing achieves the best performance (PSNR = 19.123), due to its adaptive optimization strategy for handling occlusions and environmental variations in outdoor scenes. The second best is 3DGS-MCMC Kheradmand et al. (2024) (PSNR = 18.239) on background-focused scenes with IDT Xiao et al. (2022) and VGGT, which shows robustness to initialization.
Table 3: Comparison of the performance of different raindrop removel methods. Uformer Wang et al. (2022) achieves the best performance, Restormer Zamir et al. (2022) attains performance that closely matches Uformer’s.
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
IDT Uformer Restormer IDT Uformer Restormer IDT Uformer Restormer
22.025 26.731 26.249 0.623 0.808 0.784 0.251 0.162 0.171
3DGS 3DGS-MCMC GS-W
WeatherGS
Focus Metrics Uformer Restormer IDT Uformer Restormer IDT Uformer Restormer IDT /
RD-
focused
[…]
BG-
focused
PSNR $\uparrow$ 17.906 17.741 18.094 18.219 18.148 18.239 19.123 19.074 17.818 17.124
Table 5: Comparison of COLMAP and VGGT on RaindropGS background-focused datasets with Uformer. Owing to unsuccessful camera pose estimation by COLMAP, 3DGS methods could not be applied to the raindrop-focused datasets.
3DGS 3DGS-MCMC GS-W WeatherGS
COLMAP VGGT COLMAP VGGT COLMAP VGGT COLMAP VGGT
PSNR $\uparrow$ 19.512 18.167 17.919 18.248 18.339 20.033 18.094 17.099
SSIM $\uparrow$ 0.603 0.473 0.492 0.474 0.588 0.613 0.504 0.428
LPIPS $\downarrow$ 0.384 0.467 0.454 0.489 0.562 0.433 0.400 0.447
@@ 4.3 Qualitative Comparison
Figures 3 and 4 show the qualitative results of 3DGS Kerbl et al. (2023), WeatherGS Qian et al. (2024), GS-W Zhang et al. (2024), and 3DGS-MCMC Kheradmand et al. (2024), along with their Uformer Wang et al. (2022) and Restormer Zamir et al. (2022) outputs. Scenes with background-focused images exhibit generally good performance, while raindrop-focused images pose significant reconstruction challenges due to the loss of background details and the presence of undetected raindrops. Uformer and Restormer outperform WeatherGS in restoring raindrop-degraded images. Among 3DGS variants, 3DGS suffers from detail loss, GS-W introduces artifacts, and 3DGS-MCMC offers improved quality over 3DGS. Weather-GS exhibits considerable blurriness due to limitations in handling raindrop-degraded images but shows strong multi-view consistency. In scenes that both COLMAP and VGGT successfully reconstruct, COLMAP recovers finer geometric and photometric detail. However, under raindrop interference, COLMAP fails to reconstruct any raindrop-focused scenes from its estimated camera poses and point clouds. By contrast, VGGT demonstrates superior robustness.
@@ 4.4 Discussion
Raindrop-focused and background-focused images perform distinctly in synthetic benchmarks. Raindrop-focused images suffer significant reconstruction quality loss due to the absence of background detail, highlighting deficiencies in current raindrop feature modeling and background separation. These results reveal a substantial domain gap between synthetic environments and real-world raindrop conditions, suggesting the need for future frameworks that integrate raindrop scattering models and adaptive information completion to improve performance in real precipitation scenarios, which may also facilitates polarization inference and polarization-related applicationsYuan et al. (2025c; b; a). In camera pose estimation and point cloud initialization, more robust models will stabilize reconstruction under raindrop conditions, as current methods struggle with point cloud initialization when images are focused on raindrops. 3DGS methods that can effectively handle occlusions and illumination variations will have a clear advantage. Additionally, developing better training strategies for point cloud initialization is a key area for future research.
@@ 5 Conclusion
In summary, RaindropGS offers a novel benchmark for evaluating 3DGS methods under real-world raindrop conditions. By addressing the limitations of previous synthetic datasets, we provide a more accurate representation of 3DGS performance in practical, unconstrained environments. Through the evaluation of multiple 3DGS variants, we identify the accumulated errors through camera pose estimation, point cloud initialization, raindrop removal, and 3DGS methods. Our findings highlight the strengths and weaknesses of existing approaches, offering insights into their performance under raindrop-corrupted conditions. These results underscore the need for more robust techniques to handle diverse raindrop characteristics and multi-view inconsistencies. RaindropGS not only contributes to the advancement of 3D reconstruction under challenging conditions but also lays the foundation for future research aimed at improving 3DGS performance in real-world applications.
Appendix A Appendix
A.1 Dataset Overview
Table 6 presents the statistics of our dataset, detailing the number of images under different conditions (raindrop-focused, background-focused, and ground truth) for each scene. In total, we captured 1,018 images across 11 distinct scenes.
A.2 Technical Details
Our dataset was collected using the video recording function of a fixed camera setup. The camera and glass plate were kept stationary, while the water droplet spray was adjusted to vary the density, position, and shape of the raindrops. Additionally, we modified the angle of the glass plate relative to the vertical plane to mitigate potential reflective effects.
All images were initially captured at 2K resolution. For evaluation, we uniformly downsampled the images to 1024 * 576 resolution. For single-frame raindrop removal using Uformer Wang et al. (2022) and Restormer Zamir et al. (2022), we apply full-resolution stitching to preserve spatial resolution. Similarly, during 3DGS reconstruction, we maintained the same resolution to ensure high-fidelity geometric consistency.
3DGS Kerbl et al. (2023), WeatherGS Qian et al. (2024), and 3DGS-MCMC Kheradmand et al. (2024) share similar reconstruction and rendering times, each requiring approximately 15–20 minutes on our experimental setup, with training conducted over 30,000 iterations. In contrast, GS-W Zhang et al. (2024) is officially recommended to be trained for 70,000 iterations, with total training and rendering times ranging from 150 to 180 minutes. Although GS-W delivers strong performance, its main limitation lies in its significantly longer training time.
A.3 Further Experimental Results
We conduct a thorough comparison of the 3DGS pipeline across different scenes. Table 8 summarizes the quantitative performance of 3DGS Kerbl et al. (2023), WeatherGS Qian et al. (2024), GS-W Zhang et al. (2024), and 3DGS-MCMC Kheradmand et al. (2024). On the raindrop-focused dataset, 3DGS-MCMC Kheradmand et al. (2024) achieves the highest performance on the rustydesk scene, with a PSNR of 18.572, followed closely by GS-W Zhang et al. (2024) with a PSNR of 18.463. On the background-focused dataset, GS-W Zhang et al. (2024) demonstrates the best performance on the same scene, reaching a PSNR of 21.688. Qualitative analysis in Figure 6 also shows this conclusion. These results suggest that different types of raindrop occlusions may require distinct processing strategies for optimal performance.
Table 6: Summary of image counts in our dataset across different capture conditions (raindrop-focused, background-focused, and ground truth) for each of the 11 scenes.
Scene Raindrop-focused Background-focused GT
corner 32 32 32
beartoy 29 29 29
bicycle 31 31 31
dustbin 30 30 30
flover 24 24 24
parkbear 28 28 28
popmart 31 31 31
rustdesk 28 28 28
siyuanstone 53 53 53
train 0 32 32
yingjitongdao 32 32 32
Total 318 350 350
In Table 9, we compare two of three state-of-the-art single-image raindrop removal algorithms: Uformer Wang et al. (2022) and Restormer Zamir et al. (2022). Uformer Wang et al. (2022) demonstrates slightly better performance across the evaluated metrics. In the qualitative analysis shown in Figure 5, it can be seen that Uformer Wang et al. (2022) and Restormer Zamir et al. (2022) have similar raindrop removal capabilities, but neither can completely remove raindrops. On the raindrop-focused dataset, they also introduce additional artifacts.
Table 7: Quantitative analysis of the impact of camera pose estimation and point cloud initialization using COLMAP Schonberger & Frahm (2016b) and VGGT Wang et al. (2025) on 3DGS raindrop reconstruction performance. COLMAP failed entirely on the raindrop-focused dataset, and thus only results on the background-focused dataset are reported in the table. Diagonal entries in the table indicate reconstruction failures due to unsuccessful camera pose estimation. Notably, in the cases where COLMAP succeeded, it outperformed VGGT.
VGGT (3DGS) COLMAP (3DGS) VGGT (GS‑W) COLMAP (GS‑W) VGGT (WeatherGS) COLMAP (WeatherGS)
Scene PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
BG-focused corner 17.458 0.453 0.453 / / / 19.253 0.561 0.387 / / / 17.282 0.392 0.413 / / /
dustbin 16.780 0.374 0.568 19.414 0.516 0.431 19.205 0.484 0.691 / / / 17.069 0.347 0.467 / / /
flover 14.272 0.188 0.578 16.649 0.336 0.464 14.814 0.283 0.667 / / / 9.977 0.098 0.723 10.710 0.118 0.726
parkbear 17.479 0.435 0.464 / / / 18.953 0.507 0.490 / / / 17.271 0.376 0.441 / / /
siyuanstone 20.538 0.552 0.402 20.128 0.659 0.349 21.688 0.603 0.439 21.620 0.627 0.489 21.043 0.564 0.307 19.990 0.586 0.334
yingjityongdao 16.957 0.582 0.406 / / / 18.986 0.650 0.431 / / / 16.997 0.519 0.380 / / /
Figure 5: Qualitative comparison of single-image rain removal algorithms for 3DGS raindrop reconstruction. There is no significant difference between Uformer Wang et al. (2022) and Restormer Zamir et al. (2022).
In Table 7, we compare the impact of COLMAP Schonberger & Frahm (2016b) and VGGT Wang et al. (2025) on the performance of 3DGS reconstruction in the presence of raindrop degradation, using Uformer Wang et al. (2022) for single-image restoration. COLMAP Schonberger & Frahm (2016b) achieves superior performance when it successfully completes point cloud initialization and camera pose estimation. However, COLMAP Schonberger & Frahm (2016b) fails to perform these steps across the entire raindrop-focused dataset. Furthermore, several scenes in the background-focused dataset also encounter initialization failures, resulting in the breakdown of subsequent 3DGS methods reconstruction. In the qualitative analysis shown in Figure 7, COLMAP Schonberger & Frahm (2016b) successfully provided initialization information for 3DGS raindrop reconstruction, resulting in more detailed information.
Figure 6: Qualitative comparative analysis of four 3DGS methods. Background-focused dataset are superior to raindrop-focused dataset because they provide more image details and higher reconstruction quality. WeatherGS Qian et al. (2024) has the best restoration quality for “blurred” images, while GS-W Zhang et al. (2024) performs best in overall restoration quality.
Figure 7: Qualitative analysis of the performance of COLMAP Schonberger & Frahm (2016b) and VGGT Wang et al. (2025) in 3DGS raindrop reconstruction. In successful cases, the use of COLMAP Schonberger & Frahm (2016b) enabled 3DGS methods to obtain more details.
Table 8: Quantitative comparison of the four methods: 3DGS Kerbl et al. (2023), WeatherGS Qian et al. (2024), GS-W Zhang et al. (2024), and 3DGS-MCMC Kheradmand et al. (2024). 3DGS-MCMC Kheradmand et al. (2024) achieves the best performance on the raindrop-focused dataset, while GS-W Zhang et al. (2024) performs best on the background-focused dataset.
3DGS WeatherGS GS-W 3DGS-MCMC
Scene PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
RD-focused corner 13.420 0.341 0.650 13.642 0.336 0.639 15.594 0.535 0.903 13.839 0.361 0.663
beartoy 11.616 0.415 0.629 10.929 0.365 0.643 14.053 0.607 0.729 13.525 0.499 0.599
bicycle 13.586 0.259 0.683 13.583 0.252 0.675 15.755 0.429 0.897 14.894 0.287 0.703
dustbin 15.770 0.325 0.670 14.744 0.267 0.690 18.032 0.472 0.800 17.818 0.359 0.626
flover 13.121 0.145 0.706 9.977 0.098 0.723 13.247 0.269 0.847 12.815 0.153 0.715
parkbear 14.301 0.306 0.649 12.987 0.267 0.655 15.838 0.475 0.809 14.843 0.330 0.662
popmart 13.271 0.520 0.593 12.234 0.461 0.609 15.691 0.693 0.648 15.276 0.602 0.562
rustydesk 16.927 0.367 0.659 16.209 0.351 0.651 18.463 0.514 0.811 18.572 0.385 0.665
siyuanstone 12.665 0.367 0.702 12.647 0.340 0.675 16.620 0.540 0.891 13.473 0.387 0.722
yingjityongdao 14.261 0.416 0.624 13.748 0.340 0.617 17.701 0.589 0.744 16.039 0.461 0.625
BG-focused corner 17.458 0.453 0.453 17.282 0.392 0.413 19.253 0.561 0.387 17.928 0.470 0.465
beartoy 17.249 0.625 0.435 15.611 0.584 0.442 19.474 0.711 0.415 18.888 0.662 0.421
bicycle 19.670 0.456 0.353 19.278 0.404 0.328 19.722 0.496 0.361 18.777 0.428 0.436
dustbin 16.780 0.374 0.568 17.069 0.347 0.467 19.205 0.484 0.691 19.449 0.431 0.483
flover 14.272 0.188 0.578 9.977 0.098 0.723 14.814 0.283 0.667 12.868 0.153 0.654
parkbear 17.479 0.435 0.464 17.271 0.376 0.441 18.953 0.507 0.490 17.609 0.417 0.522
popmart 17.404 0.631 0.506 15.995 0.582 0.496 17.632 0.714 0.500 17.894 0.683 0.457
rustydesk 21.255 0.484 0.429 20.721 0.419 0.365 21.498 0.543 0.499 19.695 0.423 0.564
siyuanstone 20.538 0.552 0.402 21.043 0.564 0.307 21.688 0.603 0.439 20.482 0.545 0.429
yingjityongdao 16.957 0.582 0.406 16.997 0.519 0.380 18.986 0.650 0.431 18.598 0.603 0.427
Table 9: Quantitative analysis of the performance of two single-image restoration methods, Uformer Wang et al. (2022) and Restormer Zamir et al. (2022), applied to 3DGS methods. Uformer Wang et al. (2022) achieved slightly better performance than Restormer Zamir et al. (2022) in enhancing image quality for 3D reconstruction.
Uform. (3DGS) Restorm. (3DGS) Uform. (GS‑W) Restorm. (GS‑W) Uform. (GS‑MCMC) Restorm. (GS‑MCMC)
Scene PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
RD-focused corner 13.420 0.341 0.650 13.170 0.332 0.653 15.594 0.535 0.903 15.673 0.532 0.890 13.839 0.361 0.663 13.709 0.359 0.663
beartoy 11.616 0.415 0.629 11.535 0.428 0.623 14.053 0.607 0.729 14.103 0.597 0.743 13.525 0.499 0.599 12.976 0.484 0.605
bicycle 13.586 0.259 0.683 13.469 0.248 0.671 15.755 0.429 0.897 15.314 0.428 0.904 14.894 0.287 0.703 14.755 0.283 0.686
dustbin 15.770 0.325 0.670 15.673 0.319 0.667 18.032 0.472 0.800 18.120 0.472 0.793 17.818 0.359 0.626 17.566 0.357 0.617
flover 13.121 0.145 0.706 12.972 0.143 0.712 14.814 0.283 0.667 7.013 0.002 1.066 12.815 0.153 0.715 12.708 0.152 0.715
parkbear 14.301 0.306 0.649 14.246 0.301 0.648 15.838 0.475 0.809 16.013 0.479 0.813 14.843 0.330 0.662 15.176 0.336 0.653
popmart 13.271 0.520 0.593 13.535 0.536 0.585 15.691 0.693 0.648 16.119 0.695 0.647 15.276 0.602 0.562 15.095 0.605 0.559
rustydesk 16.927 0.367 0.659 17.251 0.367 0.655 18.463 0.514 0.811 18.219 0.515 0.800 18.572 0.385 0.665 18.373 0.382 0.662
siyuanstone 12.665 0.367 0.702 12.707 0.368 0.697 19.620 0.540 0.891 15.997 0.533 0.895 13.473 0.387 0.722 13.732 0.385 0.711
yingjityongdao 14.261 0.416 0.624 17.379 0.588 0.391 17.701 0.589 0.744 17.425 0.585 0.733 16.039 0.461 0.625 15.957 0.456 0.621
BG-focused corner 17.458 0.453 0.453 17.570 0.450 0.442 19.253 0.561 0.387 19.101 0.558 0.366 17.928 0.470 0.465 17.944 0.477 0.452
beartoy 17.249 0.625 0.435 17.128 0.609 0.449 19.474 0.711 0.415 19.755 0.701 0.426 18.888 0.662 0.421 18.877 0.651 0.430
bicycle 19.670 0.456 0.353 19.379 0.447 0.373 19.722 0.496 0.361 19.412 0.486 0.402 18.777 0.428 0.436 18.615 0.425 0.451
dustbin 16.780 0.374 0.568 16.258 0.358 0.548 19.205 0.484 0.691 19.701 0.487 0.651 19.449 0.431 0.483 19.194 0.428 0.459
flover 14.272 0.188 0.578 14.154 0.175 0.582 14.814 0.283 0.667 14.362 0.275 0.686 12.868 0.153 0.654 12.832 0.151 0.649
parkbear 17.479 0.435 0.464 17.513 0.431 0.449 18.953 0.507 0.490 18.931 0.504 0.473 17.609 0.417 0.522 17.507 0.407 0.520
popmart 17.404 0.631 0.506 16.351 0.611 0.513 17.632 0.714 0.500 17.451 0.709 0.518 17.894 0.683 0.457 18.645 0.682 0.452
rustydesk 21.255 0.484 0.429 20.854 0.476 0.402 21.498 0.543 0.499 21.345 0.532 0.395 19.695 0.423 0.564 19.316 0.414 0.551
siyuanstone 20.538 0.552 0.402 20.824 0.546 0.397 21.688 0.603 0.439 21.572 0.597 0.442 20.482 0.545 0.429 20.179 0.538 0.438
yingjityongdao 16.957 0.582 0.406 17.379 0.588 0.391 18.986 0.650 0.431 19.115 0.639 0.432 18.598 0.603 0.427 18.377 0.601 0.423

### v2

@@ 3.3 Raindrop-aware 3DGS Evaluation
With the estimated camera poses and initialized point cloud, we proceed to evaluate four representative 3DGS methods: 3DGS [kerbl20233d], WeatherGS [qian2024weathergs], GS-W [zhang2024GS-W], and 3DGS-MCMC [kheradmand2024MCMCGS]. Among these, 3DGS [kerbl20233d] serves as a standard baseline for 3D Gaussian splatting. WeatherGS [qian2024weathergs] incorporates single-image raindrop removal, so we omit the explicit raindrop removal step in its data processing pipeline. GS-W [zhang2024GS-W] is specifically designed for challenging conditions and unconstrained image collections, making it more robust to inconsistent multi-view inputs. 3DGS-MCMC [kheradmand2024MCMCGS], on the other hand, does not rely on accurate point cloud initialization. Each of the aforementioned methods has its own advantages, making their evaluation in our benchmark both meaningful and insightful.
Table 3: Comparison of the performance of different raindrop removal and restoration methods. BG-focused = background-focused, RD-focused = raindrop-focused. All model weights are taken from Raindrop Clarity and evaluated under the officially recommended settings. We highlight the best and second-best results for both background-focused and raindrop-focused datasets.
AtGAN IDT Restormer Uformer
Metric BG-focused RD-focused BG-focused RD-focused BG-focused RD-focused BG-focused RD-focused
PSNR $\uparrow$ 22.366 19.758 24.203 19.847 28.442 24.055 28.997 24.465
SSIM $\uparrow$ 0.683 0.458 0.765 0.481 0.861 0.708 0.880 0.736
LPIPS $\downarrow$ 0.233 0.365 0.157 0.344 0.108 0.235 0.106 0.218
Figure 4: Qualitative Comparison across data pre-processing: Uformer and Restormer exhibit comparable raindrop removal capability. While COLMAP, when successful in estimating camera poses, reconstructs more fine-grained details than VGGT.
@@ 4.1 Implementation Details
We standardize all scene images to a resolution of 1024 * 576 for uniform comparison. During VGGT-based camera pose estimation and point cloud initialization, we follow the preprocessing protocol of VGGT [wang2025vggt], resizing each input image to 518 * 518 before processing. Likewise, for Restormer [zamir2022restormer] and IDT [xiao2022IDT], we downscale images to 128 * 128; for Uformer [wang2022uformer] and AtGAN [qian2018atgan], we downscale images to 256 * 256, and then tile the results to restore the original resolution.
For all 3DGS methods, we follow the official optimization settings: 30,000 iterations for 3DGS, WeatherGS, and 3DGS-MCMC, and 70,000 iterations for GS-W.
For 3DGS-MCMC, we set the initial point cloud size to 100,000, based on the number of points that VGGT can initialize in our dataset. All models are implemented in PyTorch and trained on 8 NVIDIA RTX 3090 GPUs.
@@ 4.2 Quantitative Comparison
Table 2 compares the impact of background-focused (BG-focused) and raindrop-focused (RD-focused) captures on 3DGS performance using VGGT [wang2025vggt]. For the original 3DGS method, raindrop-focused images exhibit a 4 dB drop compared to background-focused images, due to background blur and light refraction. VGGT processes all scenes but generates a point cloud with 0 points for raindrop-focused images, for which we use a randomly initialized point cloud.
Table 5 and Table 4 compares the performance of VGGT [wang2025vggt] and COLMAP [schonberger2016colmap]. For camera pose estimation, we use VGGT and COLMAP to estimate poses on the ground-truth images and compare these estimates with the poses obtained from the corresponding images. VGGT yields more accurate camera pose estimates; both methods accurately recover poses for background-focused images but exhibit substantial performance degradation on raindrop-focused images. In terms of point cloud initialization, VGGT outperforms COLMAP for background point clouds, yet it fails to initialize the raindrop-focused dataset.
Table 3 reports the performance of different deraining and restoration methods on raindrop-affected images. Uformer [wang2022uformer] and Restormer [zamir2022restormer] show comparable results, with Uformer [wang2022uformer] slightly outperforming Restormer [zamir2022restormer] in reconstruction metrics, while IDT [xiao2022IDT] lags substantially behind the other two methods.
Table 4: Quantitative analysis of the impact of camera pose estimation and point cloud initialization using COLMAP [schonberger2016colmap] and VGGT [wang2025vggt] on 3DGS raindrop reconstruction performance. Three reconstruction methods, 3DGS [kerbl20233d], GS-W [zhang2024GS-W], and WeatherGS [qian2024weathergs], are used for testing. COLMAP failed entirely on the raindrop-focused dataset; only background-focused results are reported. $\times$” indicates reconstruction failures caused by unsuccessful camera pose estimation. We highlight the best and second-best results for each scene.
VGGT (3DGS) COLMAP (3DGS) VGGT (GS-W) COLMAP (GS-W) VGGT (WeatherGS) COLMAP (WeatherGS)
Scene PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
corner 17.458 0.453 0.453 $\times$ $\times$ $\times$ 19.253 0.561 0.387 $\times$ $\times$ $\times$ 17.282 0.392 0.413 $\times$ $\times$ $\times$
dustbin 16.780 0.374 0.568 19.414 0.516 0.431 19.205 0.484 0.691 $\times$ $\times$ $\times$ 17.069 0.347 0.467 $\times$ $\times$ $\times$
flower 14.272 0.188 0.578 16.649 0.336 0.464 14.814 0.283 0.667 $\times$ $\times$ $\times$ 9.977 0.098 0.723 10.710 0.118 0.726
parkbear 17.479 0.435 0.464 $\times$ $\times$ $\times$ 18.953 0.507 0.490 $\times$ $\times$ $\times$ 17.271 0.376 0.441 $\times$ $\times$ $\times$
stone 20.538 0.552 0.402 20.128 0.659 0.349 21.688 0.603 0.439 21.620 0.627 0.489 21.043 0.564 0.307 19.990 0.586 0.334
door 16.957 0.582 0.406 $\times$ $\times$ $\times$ 18.986 0.650 0.431 $\times$ $\times$ $\times$ 16.997 0.519 0.380 $\times$ $\times$ $\times$
Table 5: Comparison of camera pose estimation (AUC@30 [wang2023posediffusion]) and the number of points in the point cloud between VGGT and COLMAP on background- and raindrop-focused datasets. BG-focused = background-focused, RD-focused = raindrop-focused.
@@
Num. of Points 69401.11 x 5476.89 302.50
For 3DGS methods, GS-W [zhang2024GS-W] with VGGT [wang2025vggt] and Uformer [wang2022uformer] preprocessing achieves the best performance (PSNR = 19.123), due to its adaptive optimization strategy for handling occlusions and environmental variations in outdoor scenes. The second best is 3DGS-MCMC [kheradmand2024MCMCGS] (PSNR = 18.239) on background-focused scenes with IDT [xiao2022IDT] and VGGT, which shows robustness to initialization.
@@ 4.3 Qualitative Comparison
Figures 3 and 4 show the qualitative results of 3DGS [kerbl20233d], WeatherGS [qian2024weathergs], GS-W [zhang2024GS-W], and 3DGS-MCMC [kheradmand2024MCMCGS], along with their Uformer [wang2022uformer] and Restormer [zamir2022restormer] outputs. Scenes with background-focused images exhibit generally good performance, while raindrop-focused images pose significant reconstruction challenges due to the loss of background details and the presence of undetected raindrops. Uformer and Restormer outperform WeatherGS in restoring raindrop-degraded images. Among 3DGS variants, 3DGS suffers from detail loss, GS-W introduces artifacts, and 3DGS-MCMC offers improved quality over 3DGS. Weather-GS exhibits considerable blurriness due to limitations in handling raindrop-degraded images but shows strong multi-view consistency.
In scenes that both COLMAP and VGGT successfully reconstruct, COLMAP recovers finer geometric and photometric detail. However, under raindrop interference, COLMAP fails to reconstruct any raindrop-focused scenes from its estimated camera poses and point clouds. By contrast, VGGT demonstrates superior robustness.
@@ 4.4 Discussion
Our benchmark reveals how raindrop-induced degradations affect the entire 3DGS reconstruction pipeline. We summarize three key insights as follows.
Impact of Different Focus Conditions As shown in Table 5, raindrop-affected images significantly degrade the accuracy of camera pose estimation and point cloud initialization, especially under raindrop-focused conditions. Although VGGT provides improved robustness on background-focused images, it still fails easily when strong blur and focal-plane shifts occur in raindrop-focused views. These observations suggest that future methods must explicitly handle the geometric ambiguities introduced by focus shifts and adherent raindrops. Moreover, downstream 3D reconstruction should incorporate strategies that can tolerate inaccurate poses and sparse or noisy point clouds, as these errors propagate throughout the entire 3DGS optimization process.
Impact of Raindrop Removal Methods Table 2 shows that applying robust raindrop removal methods can improve 3D reconstruction quality. However, a substantial performance gap remains, indicating that raindrop removal alone is insufficient for reliable reconstruction. In practice, downstream 3DGS methods must remain effective even when the removal results are imperfect, making it necessary to leverage multi-view information to compensate for incomplete or inaccurate restoration. Moreover, the removal quality varies across views, introducing multi-view inconsistencies that propagate into 3DGS optimization and often manifest as floaters or structural artifacts. Addressing these inconsistencies and improving robustness to imperfect raindrop removal will be important directions for future research.
Impact of 3DGS Methods Table 2 indicates the performance of various 3DGS approaches under raindrop-affected scenes. Overall, GS-W [zhang2024GS-W] achieves the best reconstruction quality, largely due to its robustness to inconsistent or transient objects. In our setting, raindrops naturally behave as view-inconsistent elements, and GS-W’s design allows it to better tolerate such variability. This observation suggests that future research should explore reconstruction strategies that explicitly account for view inconsistency, transient artifacts, or incomplete observations, enabling 3DGS pipelines to remain stable even under severe multi-view degradations.
@@ 5 Conclusion
In summary, RaindropGS offers a novel benchmark for evaluating 3DGS methods under real-world raindrop conditions. By addressing the limitations of previous synthetic datasets, we provide a more effective assessment of 3DGS performance in practical, unconstrained environments. Through the evaluation of multiple 3DGS variants, we identify the accumulated errors in camera pose estimation, point cloud initialization, raindrop removal, and 3DGS methods. Our findings highlight the strengths and weaknesses of existing approaches, offering insights into their performance under raindrop-corrupted conditions. These results underscore the need for more robust techniques to handle diverse raindrop characteristics and multi-view inconsistencies. RaindropGS not only contributes to the advancement of 3D reconstruction under challenging conditions but also lays the foundation for future research aimed at improving 3DGS performance in real-world applications.